-4x-9-4x=39. What is X ? Thank you

Found 2 solutions by rothauserc, josgarithmetic:
Answer by rothauserc(4718)

-4x -9 -4x = 39
:
add 9 to both sides of the =
:
-4x -4x = 48
:
-8x = 48
:
divide both side of = by -8
:
x = -6
:
check the answer by subsittuting for x in original equation
:
-4(-6) -9 -4(-6) = 39
:
24 -9 +24 = 39
:
48 -9 = 39
:
39 = 39
:
answer checks
:

Answer by josgarithmetic(39617)

X is a different variable. The only variable in the given equation is x.